Myers Business Systems is evaluating the introduction of a new product. The possible levels of unit sales and the probabilities of their occurrence are given next:

Possible Market Reaction Sales in Units Probabilities
Low response 35 0.20
Moderate response 45 0.20
High response 60 0.30
Very high response 70 0.30


a. What is the expected value of unit sales for the new product? (Do not round intermediate calculations and round your answer to the nearest whole unit.)

Expected value units

b. What is the standard deviation of unit sales? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Standard deviation units
Answer #1

Expected value=Respective sales*Respective probability

=(0.2*35)+(0.2*45)+(0.3*60)+(0.3*70)

=55 units

Probability Sales Probability*(Sales-Expected value)^2
0.2 35 0.2*(35-55)^2=80
0.2 45 0.2*(45-55)^2=20
0.3 60 0.3*(60-55)^2=7.5
0.3 70 0.3*(70-55)^2=67.5
Total=175

Standard deviation=[Total Probability*(Sales-Expected value)^2/Total probability]^(1/2)

=(175)^(1/2)

=13.23 units(Approx).